examine the two lines given by 2x-8y=7 and -x+4y-1=0 to determine if they are parallel, perpendicular, or neither.
Answer by TimothyLamb(4379)

---
2x - 8y = 7
-x + 4y - 1 = 0
---
put the system of linear equations into standard form
---
2x - 8y = 7
-x + 4y = 1
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
answer1:
this linear system is singular, it has no solution:
the lines are parallel
---
another method:
convert the lines to slope-intercept form:
---
2x - 8y = 7
-8y = -2x + 7
y = -2x/(-8) + 7/(-8)
y = (1/4)x - 7/8
---
-x + 4y = 1
4y = x + 1
y = (1/4)x + 1/4
---
answer2:
both lines have the same slope (1/4) so the lines are parallel
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php